OpenStudy (anonymous):

Use implicit differentiation to find the equation of the tangent line to the curve xy 3 +xy=6 at the point (3,1) . The equation of this tangent line can be written in the form y=mx+b where m=? and b=?

OpenStudy (amistre64):

is this your equation? xy^3 + xy = 6

OpenStudy (anonymous):

yes, sorry

OpenStudy (amistre64):

x'.y^3 + 3xy^2.y' + x'.y + x.y' = 0 ; where x' = 1; solve for y'

OpenStudy (amistre64):

\[y^3 + 3xy^2.y' + y + x.y' = 0\] \[3xy^2.y' + x.y' = -y^3-y\] \[y'(3xy^2 + x) = -y^3-y\] \[y' = \frac{-y^3-y}{3xy^2 + x}\]

OpenStudy (amistre64):

your point gives you your x and y value; so plug those in for your slope

OpenStudy (amistre64):

\[y' = \frac{-1-1}{3(3)(1)+3}\] \[y' = -\frac{2}{12}=-\frac{1}{6}\]

OpenStudy (amistre64):

\[tangent=(-1/6)x -(-(3)/6) + 1\] \[tangent=(-1/6)x +1/2 + 1\] \[tangent=(-1/6)x +1.5\]

OpenStudy (anonymous):

so with those you just redistribute and solve for y'?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

implicit diff is the same as explicit; in fact explicit is the case where you only have a single y term

OpenStudy (anonymous):

which makes the latter easeir :/

OpenStudy (amistre64):

\[\frac{d(y)}{dx}=\frac{dy}{dx}\] \[\frac{d(y^2)}{dx}=\frac{dy}{dx}2y\] \[\frac{d(xy)}{dx}=\frac{dx}{dx}y + x\frac{dy}{dx}\]

OpenStudy (amistre64):

its not that its easier; its just that in case of an "implicit" the solution to dy/dx involves terms of y :) when youre used to there not beig any

OpenStudy (amistre64):

the power rule is stillthe power rule; the quotitent rule is still the quotient rule; the chain rule is still the chain rule etc etc etc

OpenStudy (anonymous):

yeah, but it also makes it that much harder to shortcut, using a calculator for example.

OpenStudy (amistre64):

the only reason youre used to throwing out the derived bits; is becasue dx/dx = 1; but if an equation is derived with respect to an other variable like time: \[x^2 + xy = 67yx^2\] \[Dt(x^2 + xy = 67yx^2)\] \[\frac{dx}{dt}2x + \frac{dx}{dt}y + x\frac{dy}{dt} = \frac{dx}{dt}134yx + 67x^2\frac{dy}{dt}\]

OpenStudy (amistre64):

lol .... yeah, calcs hate multivariables